# Accuracy of the Ensemble Kalman Filter in the Near-Linear Setting

### v1

@@ 2.3 Error Estimate: Mean Field Ensemble Kalman Filter
Theorem 2.1 shows that the mean field ensemble Kalman filter error can be made arbitrarily small if the true filter is arbitrarily close to its Gaussian projection, in state-observation space. This “closeness to Gaussian” assumption can be satisfied in our setting of unbounded vector fields by considering small perturbations of affine vector fields, stated and proved in Proposition 2.2. Combining Proposition 2.2 with Theorem 2.1 gives an error estimate for the mean field ensemble Kalman filter, yielding Theorem 2.3.
@@
In what follows $(\mu^{0}_{j})_{j\in\llbracket 0,J\rrbracket}$ and $(\mu_{j})_{j\in\llbracket 0,J\rrbracket}$ denote the true filtering distributions associated with functions $(\Psi_{0},h_{0})$ and $(\Psi,h)$, respectively, initialized at the same Gaussian measure $\mathcal{N}(m_{0},C_{0})$. Furthermore, we let $\mathsf{P}_{0}$ and $\mathsf{Q}_{0}$ denote the kernel integral operators (1.7) and (1.9) defined by the specific vector fields $(\Psi_{0},h_{0}).$ By Lemma B.3, the filtering distributions have bounded second moments. Let
@@
Assume that the probability measures $(\mu_{j})_{j\in\llbracket 0,J\rrbracket}$ and $(\mu^{\rm EK}_{j})_{j\in\llbracket 0,J\rrbracket}$ are obtained respectively from the dynamical systems (1.11) and (2.4) initialized at the same Gaussian probability measure $\mu_{0}=\mu_{0}^{\rm EK}\in~{}\mathcal{G}(^{d_{u}})$. That is,
@@
Suppose that the data $Y^{\dagger}=\{y_{j}^{\dagger}\}_{j=1}^{J}$ and the matrices $(\Sigma,\Gamma)$ satisfy Assumptions (H1) and (H5). Fix $\kappa_{\Psi},\kappa_{h},\ell_{h}>0$ and let the vector fields $\Psi_{0},h_{0}$ be affine functions satisfying Assumptions (V1) and (V2), respectively, while also satisfying Assumptions (H2) and (H3) with $\kappa_{\Psi},\kappa_{h}$. Then for all $\varepsilon\in[0,1]$, there exists a constant  $C=C\bigl{(}\mathcal{M}_{q}(\mu_{0}),\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\Sigma,\Gamma,J\bigr{)}$, where $q\coloneqq{\max\{3+d_{u},4+d_{y}\}}$, such that for any $(\Psi,h)\in B_{L^{\infty}}\bigl{(}(\Psi_{0},h_{0}),\varepsilon\bigr{)}$ with $h$ satisfying Assumption (H4), it holds that
@@
Since Assumption H is satisfied, it is possible to apply the result of Theorem 2.1 to deduce that there exists $C=C\bigl{(}\mathcal{M}_{\max\{3+d_{u},4+d_{y}\}}(\mu_{0}),J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\Sigma,\Gamma\bigr{)}$ such that
@@ 2.4 Error Estimate: Finite Particle Ensemble Kalman Filter
In this subsection, we combine the results from the work in [le2009large] with stability Theorem 2.1, together with approximation Theorem 2.3, to derive a quantitative error estimate between the finite particle ensemble Kalman filter and the true filter in the non-linear setting. In order to define an appropriate metric we introduce the following class of vector fields.
@@
where $u_{j+1}^{(i)}$ evolve according to the iteration in (2.5). Suppose that the data $Y^{\dagger}=\{y_{j}^{\dagger}\}_{j=1}^{J}$ and the matrices $(\Sigma,\Gamma)$ satisfy Assumptions (H1) and (H5). Assume that the vector field $h$ is linear, and let $\kappa_{h},\ell_{h}$ be positive constants such that Assumptions (H3) and (H4) are satisfied. Furthermore, let the vector field $\Psi_{0}$ be an affine function satisfying Assumption (V1) as well as Assumption (H2) with $\kappa_{\Psi}>0$. Additionally, let $\ell_{\Psi}>0$ be a constant. Then, for all $\varepsilon\in[0,1]$, it holds that for any $\Psi\in B_{L^{\infty}}\bigl{(}\Psi_{0},\varepsilon\bigr{)}$ satisfying $|\Psi|_{C^{0,1}}\leqslant\ell_{\Psi}<\infty$, there exists a constant $C=C\bigl{(}\mathcal{M}_{q}(\mu_{0}),R_{\phi},L_{\phi},\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\ell_{\Psi},\Sigma,\Gamma,J\bigr{)}$, where $q\coloneqq{\max\{3+d_{u},4+d_{y},16\cdot 3^{J-1}\}}$, such that
$\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\mu_{J}[\phi]\Bigr{|}^{2}\right)^{1/2}\leqslant C\Bigl{(}\frac{1}{\sqrt{N}}+{\varepsilon}\Bigr{)},$
for any $\phi$ satisfying Assumption P2.
@@
    The work in [le2009large] establishes a Monte Carlo error estimate, with rate of $1/\sqrt{N}$, between the empirical measure $\mu^{{\rm EK},N}$, representing the particle ensemble Kalman filter, and the measure $\mu^{\rm EK}$ describing the evolution of the mean field ensemble Kalman filter. In particular, this holds under Assumptions (H1) and (H5) on the data and covariances of the noise processes, the linearity of the vector field $h$ and Assumption (H2) on $\Psi$ with the additional assumption that $|\Psi|_{C^{0,1}}\leqslant\ell_{\Psi}<\infty$. In Lemma D.1 we mimic the proof of [le2009large, Proposition 4.4] and  [le2009large, Theorem 5.2] to gain insight into the dependence of the constant prefactor multiplying $1/\sqrt{N}$ on the parameters of the Gaussian initial condition and the number of steps $J$.
@@
where $C$ is a constant depending on $\mathcal{M}_{\max\{3+d_{u},4+d_{y}\}}(\mu_{0}),J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\Sigma,\Gamma$.
@@
where $C$ is a constant depending on $\mathcal{M}_{16\cdot 3^{J-1}}(\mu_{0}),J,R_{\phi},L_{\phi},\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma$. In view of (2.12), combining (2.13) and (2.14) yields the desired result. ∎
@@ 3 Discussion and Future Directions
In this paper we have presented the first analysis, in the setting of a filtering problem defined by nonlinear state space dynamics, quantifying the error between the empirical measure obtained by the ensemble Kalman filter and the true filtering distribution. The analysis for the EnKF outlined is based on the proof methodology developed for particle filters in [rebeschini2015can]. This work extends the initial results of [carrillo2022ensemble] to a setting of practical interest, for which dynamical models and observation operators are allowed to be unbounded and nonlinear. The work opens a new avenue for analysis in the quantitative understanding of the performance of the ensemble Kalman methodology applied to nonlinear problems. To this end, we identify the following avenues for further work.
@@
    As surveyed in [calvello2025ensemble], the ensemble Kalman filtering methodology may be used for solving inverse problems and for sampling; it is of interest to extend this analysis to the ensemble Kalman based inversion algorithms outlined in that paper.
@@ Acknowledgments
The research of PM is supported by the projects SWIDIMS (ANR-20-CE40-0022) and CONVIVIALITY (ANR-23-CE40-0003) of the French National Research Agency. The work of AMS is supported by a Department of Defense Vannevar Bush Faculty Fellowship, and by the SciAI Center, funded by the Office of Naval Research (ONR), Grant Number N00014-23-1-2729; these grants also support the work of EC. UV is partially supported by the European Research Council (ERC) under the EU Horizon 2020 programme (grant agreement No 810367), and by the Agence Nationale de la Recherche under grants ANR-21-CE40-0006 (SINEQ) and ANR-23-CE40-0027 (IPSO). The work was initiated during The Jacques-Louis Lions Lectures given by AMS at Sorbonne Université in December 2023; the authors are grateful to Albert Cohen for hosting this event.
@@ Appendix A Auxiliary Results
Let $X$ be a random variable taking values in ${}^{d_{u}}$ with finite second moment, and let $m:=\mathbb{E}[X]$. Then it holds that
$\mathbb{E}\Bigl{[}(X-m)(X-m)^{\mathsf{T}}\Bigr{]}=\mathbb{E}\bigl{[}XX^{\mathsf{T}}\bigr{]}-mm^{\mathsf{T}}$ (A.1)
@@
$\forall a\in^{d_{u}},\qquad\mathbb{E}\Bigl{[}(X-a)(X-a)^{\mathsf{T}}\Bigr{]}\succcurlyeq\mathbb{E}\Bigl{[}(X-m)(X-m)^{\mathsf{T}}\Bigr{]}.$ (A.2)
@@
It holds that
$\mathbb{E}\Bigl{[}(X-a)(X-a)^{\mathsf{T}}\Bigr{]}$ $=\mathbb{E}\Bigl{[}\bigl{(}(X-m)+(m-a)\bigr{)}\bigl{(}(X-m)+(m-a)\bigr{)}^{\mathsf{T}}\Bigr{]}$
$=\mathbb{E}\Bigl{[}(X-m)(X-m)^{\mathsf{T}}\Bigr{]}+(m-a)(m-a)^{\mathsf{T}}.$
Setting $a=0$, we obtain (A.1). The inequality (A.2) follows by noting that the second term in the last expression is positive semidefinite. ∎
@@
Denote by $g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};m,S)$ the Lebesgue density of $\mathcal{N}(m,S)$ and by $\mathcal{S}^{n}_{\alpha}$ the set of symmetric $n\times n$ matrices $M$ satisfying
$\frac{1}{\alpha}I_{n}\preccurlyeq M\preccurlyeq\alpha I_{n}.$ (A.3)
Then for all $n\in\mathbb{N}^{+}$ and $\alpha\geqslant 1$, there exists $L_{n,\alpha}>0$ such that for all parameters $(c_{1},m_{1},S_{1})\in\times^{n}\times\mathcal{S}^{n}_{\alpha}$ and $(c_{2},m_{2},S_{2})\in\times^{n}\times\mathcal{S}^{n}_{\alpha}$,
$\lVert h\rVert_{\infty}\leqslant L_{n,\alpha}\lVert h\rVert_{1},\qquad h(y)=c_{1}g(y;m_{1},S_{1})-c_{2}g(y;m_{2},S_{2}).$ (A.4)
@@
The lemma as stated may be found in [carrillo2022ensemble, Lemma 14], where a complete proof is given. ∎
@@
Let $\mathsf{P}$ and $\mathsf{Q}$ denote the operators on probability measures given respectively in (1.7) and (1.9). Let Assumptions (H2), (H3), (H4) and (H5) be satisfied and suppose that $\mu\in\mathcal{P}(^{d_{u}})$ satisfies, for some $q>0$, the moment bound
@@
We show moment bounds in Subsection B.1, we recall that on Gaussian measures the action of the conditioning map and the Kalman transport map are equivalent in Subsection B.2, and we prove stability results in Subsection B.3.

### v2

@@ 2.3 Error Estimate: Mean Field Ensemble Kalman Filter
It is possible to deduce from the result of Theorem 2.1 that the error between the true filter and the mean field ensemble Kalman filter can be arbitrarily small if the true filter is arbitrarily close to its Gaussian projection, in state-observation space. This condition on the true filter, which we refer to as “closeness to Gaussian”, can be satisfied in the setting of unbounded vector fields by considering small perturbations of affine vector fields, as we prove in Proposition 2.2. Combining Proposition 2.2 with Theorem 2.1 gives an error estimate for the mean field ensemble Kalman filter, yielding Theorem 2.3.
@@
In what follows $(\mu^{0}_{j})_{j\in\llbracket 0,J\rrbracket}$ and $(\mu_{j})_{j\in\llbracket 0,J\rrbracket}$ denote the true filtering distributions associated with functions $(\Psi_{0},h_{0})$ and $(\Psi,h)$, respectively, initialized at the same Gaussian measure $\mathcal{N}(m_{0},C_{0})$. Furthermore, we let $\mathsf{P}_{0}$ and $\mathsf{Q}_{0}$ denote the kernel integral operators (1.7a) and (1.8a) defined by the specific vector fields $(\Psi_{0},h_{0}).$ By Lemma B.3, the filtering distributions have bounded second moments. Let
@@
Assume that the probability measures $(\mu_{j})_{j\in\llbracket 0,J\rrbracket}$ and $(\mu^{{\rm EK}}_{j})_{j\in\llbracket 0,J\rrbracket}$ are given respectively by the dynamical systems (1.9) and (2.4) initialized at the Gaussian measure $\mu_{0}=\mu_{0}^{{\rm EK}}\in~{}\mathcal{G}(^{d_{u}})$. That is,
@@
Suppose that the data $Y^{\dagger}=\{y_{j}^{\dagger}\}_{j=1}^{J}$ and the matrices $(\Sigma,\Gamma)$ satisfy Assumptions (H1) and (H2). Let the vector fields $\Psi_{0},h_{0}$ be affine functions satisfying Assumptions (V1) and (V2), respectively, while also satisfying Assumptions (H3) and (H4) with some constants $\kappa_{\Psi},\kappa_{h}>0$. Additionally, let $\ell_{\Psi}>0$ be a constant. Then there exists a constant  $C=C\bigl{(}\mathcal{M}_{q}(\mu_{0}),\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\Sigma,\Gamma,J\bigr{)}$, where $q\coloneqq{2\cdot\max\{3+d_{u},4+d_{y}\}}$, such that for all $\varepsilon\in[0,1]$ and all $(\Psi,h)\in B_{L^{\infty}}\bigl{(}(\Psi_{0},h_{0}),\varepsilon\bigr{)}$ with $h$ satisfying Assumption (H5), it holds that
@@
Since Assumption H is satisfied, it is possible to apply the result of Theorem 2.1 to deduce that there exists $C=C\bigl{(}\mathcal{M}_{2\cdot\max\{3+d_{u},4+d_{y}\}}(\mu_{0}),J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\Sigma,\Gamma\bigr{)}$ such that
@@ 2.4 Error Estimate: Finite Particle Ensemble Kalman Filter
In this subsection, we combine the results from the work in [le2009large] with stability Theorem 2.1, together with approximation Theorem 2.3, to derive a quantitative error estimate between the finite particle ensemble Kalman filter and the true filter in the nonlinear setting. In order to define an appropriate metric we introduce the following class of vector fields.
@@
where $u_{j+1}^{(i)}$ evolve according to the iteration in (2.5). Suppose that the data $Y^{\dagger}=\{y_{j}^{\dagger}\}_{j=1}^{J}$ and the matrices $(\Sigma,\Gamma)$ satisfy Assumptions (H1) and (H2). Assume that the vector field $h$ is linear, and let $\kappa_{h},\ell_{h}$ be positive constants such that Assumptions (H4) and (H5) are satisfied. Furthermore, let the vector field $\Psi_{0}$ be an affine function satisfying Assumption (V1) as well as Assumption (H3) with $\kappa_{\Psi}>0$. Additionally, let $\ell_{\Psi}>0$ be a constant. Then, there exists a constant $C=C\bigl{(}\mathcal{M}_{q}(\mu_{0}),R_{\phi},L_{\phi},\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\ell_{\Psi},\Sigma,\Gamma,J\bigr{)}$, where $q\coloneqq{2\cdot\max\{3+d_{u},4+d_{y},2\cdot 4^{J}\}}$, such that for all $\varepsilon\in[0,1]$ and all $\Psi\in B_{L^{\infty}}\bigl{(}\Psi_{0},\varepsilon\bigr{)}$ satisfying $|\Psi|_{C^{0,1}}\leqslant\ell_{\Psi}<\infty$, the following bound holds for any $\phi$ satisfying Assumption P2:
$\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\mu_{J}[\phi]\Bigr{|}^{2}\right)^{1/2}\leqslant C\Bigl{(}\frac{1}{\sqrt{N}}+{\varepsilon}\Bigr{)}.$
@@
    The work in [le2009large] establishes a Monte Carlo error estimate, with rate of $1/\sqrt{N}$, between the empirical measure $\mu^{{\rm EK},N}$, representing the particle ensemble Kalman filter, and the measure $\mu^{{\rm EK}}$ describing the evolution of the mean field ensemble Kalman filter. In particular, this holds under Assumptions (H1) and (H2) on the data and covariances of the noise processes, the linearity of the vector field $h$ and Assumption (H3) on $\Psi$ with the additional assumption that $|\Psi|_{C^{0,1}}\leqslant\ell_{\Psi}<\infty$. In Lemma D.1 we provide a self-contained proof of [le2009large, Theorem 5.2] to gain insight into the dependence of the constant prefactor multiplying $1/\sqrt{N}$ on the parameters of the Gaussian initial condition and the number of steps $J$.
@@
where $C$ is a constant depending on $\mathcal{M}_{2\cdot\max\{3+d_{u},4+d_{y}\}}(\mu_{0}),J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\Sigma,\Gamma$.
@@
where $C$ is a constant depending on $\mathcal{M}_{4^{J+1}}(\mu_{0}),J,R_{\phi},L_{\phi},\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma$. In view of (2.12), combining (2.13) and (2.14) yields the desired result. ∎
@@ 3 Discussion and Future Directions
In this paper we have presented the first analysis, in the setting of a filtering problem defined by nonlinear state space dynamics, quantifying the error between the empirical measure obtained by the finite particle ensemble Kalman filter and the true filtering distribution. The analysis for the EnKF outlined is based on the proof methodology developed for particle filters in [rebeschini2015can] and developed for the mean field ensemble Kalman filter in [carrillo2022ensemble]. Our work goes beyond [carrillo2022ensemble] by considering finite particle filters and by allowing for dynamical models and observation operators that are unbounded. The work builds substantially on the new avenue for analysis of the ensemble Kalman methodology that was introduced in [carrillo2022ensemble] but leaves open numerous avenues of investigation for further work; we detail these.
@@
    As surveyed in [calvello2025ensemble], the ensemble Kalman filtering methodology may be used for solving inverse problems and for sampling; it is of interest to extend the analysis in our paper to the ensemble Kalman based inversion algorithms outlined in that paper.
@@ Acknowledgments
The research of PM is supported by the projects SWIDIMS (ANR-20-CE40-0022) and CONVIVIALITY (ANR-23-CE40-0003) of the French National Research Agency. The work of AMS is supported by a Department of Defense Vannevar Bush Faculty Fellowship, and by the SciAI Center, funded by the Office of Naval Research (ONR), Grant Number N00014-23-1-2729. The work of EC was supported by the Resnick Sustainability Institute, and also by the SciAI Center and by the Vannevar Bush Faculty Fellowship held by AMS. UV is partially supported by the European Research Council (ERC) under the EU Horizon 2020 programme (grant agreement No 810367), and by the Agence Nationale de la Recherche under grants ANR-21-CE40-0006 (SINEQ) and ANR-23-CE40-0027 (IPSO). The work was initiated during The Jacques-Louis Lions Lectures given by AMS at Sorbonne Université in December 2023; the authors are grateful to Albert Cohen for hosting this event.
[…]
@@ Appendix A Auxiliary Results
Let $X$ be a random vector in ${}^{d_{u}}$ with finite second moment. Then it holds that
$\mathbb{E}\Bigl{[}\bigl{(}X-\mathbb{E}[X]\bigr{)}\bigl{(}X-\mathbb{E}[X]\bigr{)}^{\mathsf{T}}\Bigr{]}=\mathbb{E}\bigl{[}XX^{\mathsf{T}}\bigr{]}-\mathbb{E}[X]\mathbb{E}[X]^{\mathsf{T}}$ (A.1)
@@
$\forall a\in^{d_{u}},\qquad\mathbb{E}\Bigl{[}\bigl{(}X-a\bigr{)}\bigl{(}X-a\bigr{)}^{\mathsf{T}}\Bigr{]}\succcurlyeq\mathbb{E}\Bigl{[}\bigl{(}X-\mathbb{E}[X]\bigr{)}\bigl{(}X-\mathbb{E}[X]\bigr{)}^{\mathsf{T}}\Bigr{]}.$ (A.2)
@@
The statements follow from the following equalities:
$\mathbb{E}\Bigl{[}(X-a)(X-a)^{\mathsf{T}}\Bigr{]}$ $=\mathbb{E}\Bigl{[}\Bigl{(}\bigl{(}X-\mathbb{E}[X]\bigr{)}+\bigl{(}\mathbb{E}[X]-a\bigr{)}\Bigr{)}\Bigl{(}\bigl{(}X-\mathbb{E}[X]\bigr{)}+\bigl{(}\mathbb{E}[X]-a\bigr{)}\Bigr{)}^{\mathsf{T}}\Bigr{]}$
$=\mathbb{E}\Bigl{[}\bigl{(}X-\mathbb{E}[X]\bigr{)}\bigl{(}X-\mathbb{E}[X]\bigr{)}^{\mathsf{T}}\Bigr{]}+\bigl{(}\mathbb{E}[X]-a\bigr{)}\bigl{(}\mathbb{E}[X]-a\bigr{)}^{\mathsf{T}}.\qed$
@@
Let $g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};m,S)$ be the Lebesgue density of $\mathcal{N}(m,S)$ and $\mathcal{S}^{n}_{\alpha}$ the set of symmetric $n\times n$ matrices $M$ satisfying
$\frac{1}{\tau}I_{n}\preccurlyeq M\preccurlyeq\tau I_{n}.$
Then for any $n\in\mathbb{N}^{+}$ and $\tau\geqslant 1$, there is $L_{n,\tau}>0$ such that for all parameters $(c_{1},m_{1},S_{1})\in\times^{n}\times\mathcal{S}^{n}_{\tau}$ and $(c_{2},m_{2},S_{2})\in\times^{n}\times\mathcal{S}^{n}_{\tau}$ it holds that
$\lVert h\rVert_{\infty}\leqslant L_{n,\tau}\lVert h\rVert_{1},\qquad h(y)=c_{1}g(y;m_{1},S_{1})-c_{2}g(y;m_{2},S_{2}).$
@@
The lemma as stated may be found in [carrillo2022ensemble, Lemma A.4], where a complete proof is given. ∎
@@
Let $\mathsf{P}$ and $\mathsf{Q}$ denote the operators on probability measures given respectively in (1.7a) and (1.8a). Let Assumptions (H3), (H4), (H5) and (H2) be satisfied and suppose that $\mu\in\mathcal{P}(^{d_{u}})$ satisfies, for some $q>0$, the moment bound
@@
We establish moment bounds in Subsection B.1, we recall that on Gaussian measures the action of the conditioning map and the Kalman transport map are equivalent in Subsection B.2, and we prove stability results in Subsection B.3.